OpenStudy (anonymous):

The function f(t) = 65 sin (π/5 t) +35 models the temperature of anperiodic chemical reaction where t represents time in hours. what are the maximum and minimum temperatures of the reaction, and how long does the entire cycle take?

OpenStudy (anonymous):

A) maximum:65 degrees; minimum:35 degrees; period: pi/5 hours B) maximum:100 degrees; minimum: -30 degrees; period: pi/5 hours C) maximum: 65 degrees; minimum:35 degrees;period: 10 hours D) maximum:100 degrees; minimum:-30 degrees; period:10 hours

myininaya (myininaya):

do you know the max value of y=sin(x)?

OpenStudy (anonymous):

No, I'm not sure how to find that.

myininaya (myininaya):

have you ever graphed y=sin(x) before?

OpenStudy (anonymous):

No

myininaya (myininaya):

do you know what sin(0) and sin(pi/2) and sin(pi) and sin(3pi/2) and sin(2pi) equal?

OpenStudy (anonymous):

I'm honestly just not sure where to even start with solving this question.

myininaya (myininaya):

trying to give you a place to start can you answer my last question please?

OpenStudy (anonymous):

No, I don't know that

myininaya (myininaya):

well that could be your problem and answering this because you don't know anything about the function sin(x)

myininaya (myininaya):

I don't know where to start now

myininaya (myininaya):

can you give me a place to start?

myininaya (myininaya):

like is there anything you know about trig?

myininaya (myininaya):

like do you know the unit circle?

OpenStudy (anonymous):

I know the basics of sin and things like that, but other than that, I haven't gone over trig for quite a while.

myininaya (myininaya):

like what basics?

myininaya (myininaya):

like knowing sin(0) and sin(pi/2) and so on are pretty basic

myininaya (myininaya):

sin(0)=0 sin(pi/2)=1 sin(pi)=0 sin(3pi/2)=-1 sin(2pi)=0 so this doesn't look at all familiar?

OpenStudy (anonymous):

It looks a bit familiar.

myininaya (myininaya):

 well using that information I'm able to graph the function y=sin(x) on [0,2pi]